OpenStudy (anonymous):

A^2 * b^-3 / C^-5 * A^4 ....how do i simplify the expression

OpenStudy (anonymous):

factor out a^2 from numerator and denominator

OpenStudy (anonymous):

b^-3 = 1/b^3

OpenStudy (anonymous):

c^-5 = 1/c^5

OpenStudy (anonymous):

so final answer = c^5/a^2b^3

OpenStudy (anonymous):

Thank you, thats what I got as well as I followed along.